A random sample of fifty-four ​200-meter swims has a mean time of 3.125 minutes and a standard deviation of 0.080 minutes. A 95​% confidence interval for the population mean time is (3.107,3.143).

Construct a 95​% confidence interval for the population mean time using a standard deviation of 0.05 minutes. Which confidence interval is​ wider? Explain.

The 95​% confidence interval is ( ____ , ____ ) (Round to three decimal places as​ needed.)
Answer #1

Solution :

Given that,

= 3.125

s = 0.05

n = 54

Degrees of freedom = df = n - 1 = 54 - 1 = 53

At 95% confidence level the t is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

t /2,df = t0.025,54 = 2.006

Margin of error = E = t/2,df * (s /n)

= 2.006 * (0.05 / 54)

= 0.014

The 95% confidence interval estimate of the population mean is,

- E < < + E

3.125 - 0.014 < < 3.125 + 0.014

3.111 < < 3.139

The 95​% confidence interval is (3.111, 3.139)

The confidence interval for standard deviation of 0.08 is wider because difference between interval is greater than standard deviation of 0.05